Skysong, Inc. signed a lease for an office building for a period of 12 years. Under the lease agreement, a security deposit of $7,000 is made. The deposit will be returned at the expiration of the lease with interest compounded at 7% per year.

Click here to view the factor table.

(For calculation purposes, use 5 decimal places as displayed in the factor table provided.)

What amount will Skysong, Inc. receive at the time the lease expires? (Round answer to 2 decimal places, e.g. 25.25.)

Amount at the time the lease expires

$enter a dollar amount rounded to 2 decimal places
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=7000*(1.07)^12

=7000*2.25219

=$15765.33